Find the component form of vector with initial point (-6, 2) and terminal Point (7,-3)
Answer #1

Initial point (-6,2) terminal point (7,-3) component Farm of rector v v = (7+6) i + (-3-2) i - 13 1 - 5 ¿ Ñ = (13,-5)This is the component form